OpenStudy (anonymous):

what is the geometric mean of 7 and 11

OpenStudy (kaydra14):

7/x = x/11 x = sqrt(7 x 11) = sqrt(77), about 8.775

OpenStudy (gorv):

welcome to OS

OpenStudy (gorv):

geomatric mean = (multiplication of all number )^(1/total number )

OpenStudy (gorv):

total number here= 2

OpenStudy (shinalcantara):

for geometric progression: \[a_n = a_1 r ^{n-1}\] first thing to do is get the value of r for you to get the mean let's say that \[a_3 = 11 ; a_1 = 7; n=3\] \[11 = 7 r ^{3-1}\] \[\frac{ 11 }{ 7 } = r ^{2}\] getting the square root \[r = \frac{ \sqrt{77} }{ 7 }\] since the mean in this case is just the second term then you'll have it as: \[a_2 = 7(\frac{ \sqrt{77} }{ 7 })^{2-1}\] \[a_2 = \sqrt{77}\]